# REACT3D: Recovering Articulations for Interactive Physical 3D Scenes

### v1

@@ Abstract
Interactive 3D scenes are increasingly vital for embodied intelligence, yet existing datasets remain limited due to the labor-intensive process of annotating part segmentation, kinematic types, and motion trajectories. We present REACT3D, a scalable zero-shot framework that converts static 3D scenes into simulation-ready interactive replicas with consistent geometry, enabling direct use in diverse downstream tasks. Our contributions include: (i) openable-object detection and segmentation to extract candidate movable parts from static scenes, (ii) articulation estimation that infers joint types and motion parameters, (iii) hidden-geometry completion followed by interactive object assembly, and (iv) interactive scene integration in widely supported formats to ensure compatibility with standard simulation platforms. We achieve state-of-the-art performance on detection/segmentation and articulation metrics across diverse indoor scenes, demonstrating the effectiveness of our framework and providing a practical foundation for scalable interactive scene generation, thereby lowering the barrier to large-scale research on articulated scene understanding. Our project page is react3d.github.io.
@@ III-A Preliminaries
Problem formulation
Given a static 3D scene representation $\mathbf{S}$, our goal is to construct an interactive digital twin $\hat{\mathbf{S}}$ that augments the original geometry with articulated objects while preserving the static background (Fig. 2). The input $\mathbf{S}$ consists of a point cloud or a mesh $\mathbf{P}$ with per-vertex colors, captured by a registered RGB-D sensor array, together with a set of calibrated frames $\{(\mathbf{I}_{i},\mathbf{D}_{i},\mathbf{T}_{i},\mathbf{K}_{i})\}_{i=1}^{N}$, where $\mathbf{I}_{i}$ is an RGB image, $\mathbf{D}_{i}$ its depth map, $\mathbf{T}_{i}\in SE(3)$ the extrinsic pose, and $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ the intrinsic matrix. The output $\hat{\mathbf{S}}$ is decomposed as
@@
Openable object
We define an openable object as one that contains a movable part actuated by either a prismatic or a revolute joint [21]. For a prismatic joint, the part translates along a fixed axis $\mathbf{a}$ with displacement $s\in[0,\rho]$, while for a revolute joint, it rotates about a unit axis $\mathbf{a}$ through origin $\mathbf{o}$ with angle $\theta\in[0,\rho]$. The motion of any point $\mathbf{x}$ belonging to the movable part in its closed configuration is
@@ III-B Openable object detection and segmentation
Interactive scene understanding requires semantic knowledge of which objects afford interaction (e.g., doors or drawers), beyond the appearance and geometry captured in static 3D scenes. As the first stage of our framework, openable object detection identifies candidate interactive instances, while part segmentation delineates their movable components from the surrounding structure to obtain aligned masks or meshes. This decomposes the input static scene into instance-level proposals that serve as the input domain for downstream articulation estimation, thereby determining the extent of interactivity in the resulting scene.
Open-vocabulary 2D detection and segmentation
@@
Indoor scenes include diverse openable objects not captured by standard labels, motivating an open-vocabulary detection strategy to maximize interactive coverage. We therefore first process the RGB frames $\{\mathbf{I}_{i}\}_{i=1}^{N}$ of the static scene to recognize semantic tags of objects present in the scene using RAM++ [22]. Let $\mathcal{T}=\bigcup_{i=1}^{N}\texttt{RAM++}(\mathbf{I}_{i})$ denote the aggregated tag set. We then apply a VLM-based filtering method built on LLaVA [23] to retain only tags that satisfy our definition of openable objects, yielding $\mathcal{T}_{\mathrm{open}}=\{\,t\in\mathcal{T}\mid f_{\mathrm{VLM}}(t)=1\,\}$. Finally, Grounded SAM [24] is applied with prompts $\mathcal{T}_{\mathrm{open}}$ on each frame $\mathbf{I}_{i}$ to localize and segment movable parts, producing a set of 2D instance masks $\{\mathbf{m}_{i,k}\}_{k=1}^{n_{i}}$, where $n_{i}$ denotes the number of detected movable parts in frame $\mathbf{I}_{i}$, which are post-processed with hole filling to close interior gaps, as Grounded SAM can miss door handles. Unlike prior approaches restricted to predefined labels, our open-vocabulary method mitigates label bias and yields better coverage of long-tail openable objects.
2D-to-3D segmentation via multi-view fusion
@@
Zero-shot 3D instance segmentation methods [25] operating on point clouds or meshes rarely yield reliable part-level segmentation. We therefore adopt a 2D-to-3D multi-view fusion paradigm, similar to DRAWER, to lift robust 2D masks into 3D. For each frame, we construct the model-view-projection matrix from $\mathbf{T}_{i},\mathbf{K}_{i}$ to recover the camera view. The scene mesh is then rasterized under each view, and the front-most faces covered by each mask are selected as its 3D projection. To recover the full structure of each object, as they are not always fully visible from a single view, we fuse the per-mask projections by building a face graph and applying Louvain community detection [26]. Unlike DRAWER, which retains only one mask above the Intersection over Union (IoU) threshold for each fusion result, our approach ranks the masks by IoU and preserves the top-$k$ views that exceed the threshold, providing multiple perspectives that improve the robustness of subsequent articulation estimation. In addition, we replace its VLM-based filtering of part masks, which discards candidates without visible handles, with an alternative validation step at the articulation estimation stage (Sec. III-C). This design allows the detection of openable objects even when handles are absent or visually nonsalient, as shown by the comparison in Fig. 6. We then re-segment each object by invoking SAM with point prompts seeded from previously fused 3D proposals.
@@
Another challenge in 2D-to-3D lifting is that segmentation often produces the whole object rather than its movable part, e.g., the whole cabinet instead of just the door. To address this, we refine each segmented 3D mesh by applying RANSAC [27] to estimate an approximately vertical, finite-thickness plane and retain the estimate with the largest contour area. Naive planar clipping, however, may fail to capture protruding handles. To preserve such structures while discarding static body regions, let $\mathbf{n}$ denote the normal of the estimated plane and $\mathbf{v}_{i}$ the viewing direction derived from the camera pose $\mathbf{T}_{i}$ of the frame that produced the segmentation. We choose the inward-pointing normal such that $\langle\mathbf{n},\mathbf{v}_{i}\rangle>0$, and discard the geometry behind the plane along this direction. This produces high-quality 3D movable-part meshes $M^{\mathrm{obj}}_{j}=(V_{j},F_{j})$, where $V_{j}$ and $F_{j}$ denote the vertex and face sets of the segmented mesh, together with the top-$k$ RGB views and their associated masks $\{\mathbf{m}_{j,i}\}_{i=1}^{k}$. The resulting part meshes are spatially aligned with the input static scene and provide the foundation for generating interactive objects.
@@ III-C Articulation estimation
To enable meaningful manipulation, each movable part must be endowed with kinematic parameters that define its motion type and limits. Accordingly, for every movable part, we estimate an articulation tuple $\phi(m)=\big(t,\mathbf{o},\mathbf{a},\rho\big)$. In our framework, we leverage OPDMulti [6, 8] to obtain initial articulation estimates from a single image, producing masks of movable parts together with their joint type, origin, axis, and motion range.
Candidate filtering
Since OPDMulti can detect multiple movable parts from a single RGB frame, it offers the opportunity to further validate the candidate openable objects obtained from the segmentation stage. For each segmented 3D mesh, we utilize the pre-saved top-$k$ best-view images and their SAM-generated masks $M_{\mathrm{SAM}}=\{\mathbf{m}_{i}\}_{i=1}^{k}$. We then run OPDMulti on these views and compare its predicted movable-part masks $M_{\mathrm{OPDM}}$ against $M_{\mathrm{SAM}}$. A candidate is considered valid if the IoU between the two masks exceeds a threshold in at least one of the top-$k$ views; otherwise, the 3D mesh is discarded as an invalid case.
Articulation refinement
While OPDMulti provides a reasonable initial articulation, direct application to interactive objects may fail to reproduce realistic joint behavior. To recover faithful interactions, we refine the initial estimates for each validated movable part by computing an oriented bounding box (OBB) of the part mesh with COMPAS [28]. The OBB is denoted $(\mathbf{U},\mathbf{c},\mathbf{s})$, where $\mathbf{U}=[\mathbf{u}_{1}\,\mathbf{u}_{2}\,\mathbf{u}_{3}]$ is a set of orthogonal edges, $\mathbf{c}$ is the center, and $\mathbf{s}=(s_{1},s_{2},s_{3})$ are the side lengths with $s_{1}\geq s_{2}\geq s_{3}$. Since openable parts are typically plate-like, we define the front face $P_{\mathrm{front}}$ as the OBB face spanned by the two longest edges $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$ with unit normal $\mathbf{n}_{\mathrm{front}}=\mathbf{u}_{1}\times\mathbf{u}_{2}$. Let the initial articulation be $\phi(m)=(t,\mathbf{o},\mathbf{a},\rho)$. For prismatic joints, we update only the joint axis direction by aligning it with the front-face normal,
@@ III-E Interactive Scene Integration
To ensure clean assembly of the final interactive scene, we perform a last deduplication pass. Let $\{S_{i}\}$ denote movable-part meshes with point sets $\{X_{i}\}$. We compute the IoU directly on these point sets, $\mathrm{IoU}(X_{i},X_{j})=\tfrac{|X_{i}\cap X_{j}|}{|X_{i}\cup X_{j}|}$. The procedure has two stages. Stage 1 (pairwise pruning): for every pair $(i,j)$, if $\mathrm{IoU}(X_{i},X_{j})\geq\tau_{\mathrm{dup}}$, we keep the mesh with greater information content (measured by point count) and discard the other, producing a reduced set $\{\hat{S}_{i}\}$ with point sets $\{\hat{X}_{i}\}$. Stage 2 (subdivision pruning): duplicates can also appear when a larger mesh can be represented by the union of several smaller, complementary meshes while no single pair exceeds $\tau_{\mathrm{dup}}$. We sort $\{\hat{S}_{i}\}$ by $|\hat{X}_{i}|$ in descending order and iterate from large to small so that larger parts can be explained by smaller ones first. For each $\hat{S}_{i}$, we collect candidate smaller meshes $\mathcal{C}_{i}=\{\,j:|\hat{X}_{j}|<|\hat{X}_{i}|,\ \mathrm{IoU}(\hat{X}_{i},\hat{X}_{j})\geq\tau_{\mathrm{low}}\,\}$, where $\tau_{\mathrm{low}}$ is a small IoU threshold, and enumerate their combinations. If there exists a subset $\mathcal{U}\subseteq\mathcal{C}_{i}$ such that $\mathrm{IoU}\!\left(\hat{X}_{i},\ \bigcup_{j\in\mathcal{U}}\hat{X}_{j}\right)\geq\tau_{\mathrm{dup}}$, we remove $\hat{S}_{i}$ as a redundant superset. In the end, we remove all redundant movable parts together with their associated inner boxes before assembling the final interactive scene.
Scene assembly
@@
We integrate interactive objects into the original static scene by carving the background geometry using the world-frame OBBs of deduplicated movable parts. Let the original static scene mesh be $\mathbf{P}$ and the set of movable-part OBBs be $\{B_{j}\}$. To remove residual fragments from imperfect segmentation and spurious points in the original scan, we delete all vertices lying inside any $B_{j}$ together with their incident faces. This yields a cleaned background mesh $\mathbf{P}_{\mathrm{rem}}\;=\;\mathbf{P}\setminus\bigcup_{j}B_{j}$. We then assemble the final interactive scene by combining $\mathbf{P}_{\mathrm{rem}}$ with the interactive objects, ensuring a clean integration without interpenetration.
Texture generation
@@
As both the original scene mesh and the segmented movable-part meshes provide only per-vertex colors, we generate texture maps based on InstantTexture [29]. We first perform UV unwrapping with Xatlas [30] on the movable parts and the cleaned remainder mesh. Given the UV parameterization, we rasterize each triangle in atlas space and compute per-texel colors by barycentric interpolation of the source vertex colors, yielding dense textures aligned to the unwrapped charts. We then repair uncovered texels caused by sampling gaps or occlusions via inpainting, followed by smoothing with low-pass filtering to attenuate aliasing across chart boundaries. Finally, the generated textures are bound to their corresponding meshes and, together with the refined articulation parameters, exported as .dae assets that are referenced in the scene description files.
Export and simulator support
@@
To ensure portability across visualization and simulation stacks, we export each interactive scene in two widely adopted formats: URDF for the robotics ecosystem and USD for graphics/physics engines. The exports package the cleaned background, articulated interactive objects, and textures, enabling direct deployment in PyBullet [31], ROS [32], and Isaac Sim [33]. Fig. 4 shows the simulated interactive scenes generated by REACT3D from different inputs in Isaac Sim. For ROS, we additionally provide a lightweight GUI for manipulating the interactive scene during visualization and benchmarking (see Fig. 5).
@@
Method Representation #Openable Objects $\tau^{*}=0.25$ $\tau^{*}=0.50$
@@
URDFormer Mesh 226 $0.429$ $0.226$ $0.296$ $0.361$ $0.190$ $0.249$
DRAWER (GT-mesh) Points & Mesh 540 $0.434$ $0.141$ $0.213$ $0.326$ $0.106$ $0.159$
Ours Points & Mesh 540 0.731 0.393 0.511 0.666 0.357 0.465

### v3

# Recovering Articulations for Interactive Physical 3D Scenes

@@ Abstract
Interactive 3D scenes are increasingly vital for embodied intelligence, yet existing datasets remain limited due to the labor-intensive process of annotating part segmentation, kinematic types, and motion trajectories. We present REACT3D, a scalable zero-shot framework that converts static 3D scenes into simulation-ready interactive replicas with consistent geometry, enabling direct use in diverse downstream tasks. Our contributions include: (i) openable-object detection and segmentation to extract candidate movable parts from static scenes, (ii) articulation estimation that infers joint types and motion parameters, (iii) hidden-geometry completion followed by interactive object assembly, and (iv) interactive scene integration in widely supported formats to ensure compatibility with standard simulation platforms. We achieve state-of-the-art performance on detection/segmentation and articulation metrics across diverse indoor scenes, demonstrating the effectiveness of our framework and providing a practical foundation for scalable interactive scene generation, thereby lowering the barrier to large-scale research on articulated scene understanding.
@@ III-A Preliminaries
Problem formulation.  Given a static 3D scene representation $\mathbf{S}$, our goal is to construct an interactive digital twin $\hat{\mathbf{S}}$ that augments the original geometry with articulated objects while preserving the static background (Fig. 2). The input $\mathbf{S}$ consists of a point cloud or a mesh $\mathbf{P}$ with per-vertex colors, captured by a registered RGB-D sensor array, together with a set of calibrated frames $\{(\mathbf{I}_{i},\mathbf{D}_{i},\mathbf{T}_{i},\mathbf{K}_{i})\}_{i=1}^{N}$, where $\mathbf{I}_{i}$ is an RGB image, $\mathbf{D}_{i}$ its depth map, $\mathbf{T}_{i}\in SE(3)$ the extrinsic pose, and $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ the intrinsic matrix. The output $\hat{\mathbf{S}}$ is decomposed as
@@
Openable object.  We define an openable object as one that contains a movable part actuated by either a prismatic or a revolute joint [13]. For a prismatic joint, the part translates along a fixed axis $\mathbf{a}$ with displacement $s\in[0,\rho]$, while for a revolute joint, it rotates about a unit axis $\mathbf{a}$ through origin $\mathbf{o}$ with angle $\theta\in[0,\rho]$. The motion of any point $\mathbf{x}$ belonging to the movable part in its closed configuration is
@@ III-B Openable object detection and segmentation
Interactive scene understanding requires semantic knowledge of object affordances beyond static 3D geometry. The detection module is utilized to identify candidate interactive instances, while part segmentation delineates movable components from the structure to obtain aligned meshes. This decomposes the static scene into instance-level proposals for downstream articulation estimation, determining the extent of interactivity in the resulting scene.
@@
Open-vocabulary 2D detection and segmentation.  Indoor scenes include diverse openable objects not captured by standard labels, motivating an open-vocabulary detection strategy to maximize interactive coverage. We therefore first process the RGB frames $\{\mathbf{I}_{i}\}_{i=1}^{N}$ of the static scene to recognize semantic tags of objects present in the scene using RAM++ [33]. Let $\mathcal{T}=\bigcup_{i=1}^{N}\texttt{RAM++}(\mathbf{I}_{i})$ denote the aggregated tag set. We then apply a VLM-based filtering method built on LLaVA [14] to retain only tags that satisfy our definition of openable objects, yielding $\mathcal{T}_{\mathrm{open}}=\{\,t\in\mathcal{T}\mid f_{\mathrm{VLM}}(t)=1\,\}$. Finally, Grounded SAM [24] is applied with prompts $\mathcal{T}_{\mathrm{open}}$ on each frame $\mathbf{I}_{i}$ to localize and segment movable parts, producing a set of 2D instance masks $\{\mathbf{m}_{i,k}\}_{k=1}^{n_{i}}$, where $n_{i}$ denotes the number of detected movable parts in frame $\mathbf{I}_{i}$, which are post-processed with hole filling to close interior gaps, as Grounded SAM can miss door handles. Unlike prior approaches restricted to predefined labels, our open-vocabulary method mitigates label bias and yields better coverage of long-tail openable objects.
@@
2D-to-3D segmentation via multi-view fusion.  Zero-shot 3D instance segmentation methods [25] operating on point clouds or meshes rarely yield reliable part-level segmentation. We therefore adopt a 2D-to-3D multi-view fusion paradigm, similar to DRAWER, to lift robust 2D masks into 3D. For each frame, we construct the model-view-projection matrix from $\mathbf{T}_{i},\mathbf{K}_{i}$ to recover the camera view. The scene mesh is then rasterized under each view, and the front-most faces covered by each mask are selected as its 3D projection. To recover the full structure of each object, as they are not always fully visible from a single view, we fuse the per-mask projections by building a face graph and applying Louvain community detection [5]. Unlike DRAWER, which retains only one mask above the Intersection over Union (IoU) threshold for each fusion result, our approach ranks the masks by IoU and preserves the top-$k$ views that exceed the threshold, providing multiple perspectives that improve the robustness of subsequent articulation estimation. In addition, we replace its VLM-based filtering of part masks, which discards candidates without visible handles, with an alternative validation step at the articulation estimation stage (Sec. III-C). This design allows the detection of openable objects even when handles are absent or visually nonsalient, as shown by the comparison in Fig. 3. We then re-segment each object by invoking SAM with point prompts seeded from previously fused 3D proposals.
Input
DRAWER
REACT3D
Figure 3: Qualitative comparison. Red and blue arrows denote revolute and prismatic joints. Our method outperforms DRAWER on handle-less objects (top) and correctly classifies prismatic drawers that DRAWER mistakes for revolute joints (bottom).
@@
Another challenge in 2D-to-3D lifting is that segmentation often produces the whole object rather than its movable part. To address this, we refine each segmented 3D mesh by applying RANSAC [7] to estimate an approximately vertical, finite-thickness plane and retain the estimate with the largest contour area. Naive planar clipping, however, may fail to capture protruding handles. To preserve such structures while discarding static body regions, let $\mathbf{n}$ denote the normal of the estimated plane and $\mathbf{v}_{i}$ the viewing direction derived from the camera pose $\mathbf{T}_{i}$ of the frame that produced the segmentation. We choose the inward-pointing normal such that $\langle\mathbf{n},\mathbf{v}_{i}\rangle>0$, and discard the geometry behind the plane along this direction. This produces high-quality 3D movable-part meshes $M^{\mathrm{obj}}_{j}=(V_{j},F_{j})$, where $V_{j}$ and $F_{j}$ denote the vertex and face sets of the segmented mesh, together with the top-$k$ RGB views and their associated masks $\{\mathbf{m}_{j,i}\}_{i=1}^{k}$. The resulting part meshes are spatially aligned with the input static scene and provide the foundation for generating interactive objects.
@@ III-C Articulation estimation
To enable meaningful manipulation, each movable part must be endowed with kinematic parameters that define its motion. Accordingly, for every movable part, we estimate an articulation tuple $\phi(m)=\big(t,\mathbf{o},\mathbf{a},\rho\big)$. In our framework, we leverage OPDMulti [10, 27] to obtain initial articulation estimates and corresponding movable part masks.
Candidate filtering.  Since OPDMulti can detect multiple movable parts from a single RGB frame, it offers the opportunity to further validate the candidate openable objects obtained from the segmentation stage. For each segmented 3D mesh, we utilize the pre-saved top-$k$ best-view images and their SAM-generated masks $M_{\mathrm{SAM}}=\{\mathbf{m}_{i}\}_{i=1}^{k}$. We then run OPDMulti on these views and compare its predicted movable-part masks $M_{\mathrm{OPDM}}$ against $M_{\mathrm{SAM}}$. A candidate is considered valid if the IoU between the two masks exceeds a threshold in at least one of the top-$k$ views; otherwise, the 3D mesh is discarded as an invalid case.
Articulation refinement.  While OPDMulti provides a reasonable initial articulation, direct application to interactive objects may fail to reproduce realistic joint behavior. To recover faithful interactions, we refine the initial estimates for each validated movable part by computing an oriented bounding box (OBB) of the part mesh with COMPAS [19]. The OBB is denoted $(\mathbf{U},\mathbf{c},\mathbf{s})$, where $\mathbf{U}=[\mathbf{u}_{1}\,\mathbf{u}_{2}\,\mathbf{u}_{3}]$ is a set of orthogonal edges, $\mathbf{c}$ is the center, and $\mathbf{s}=(s_{1},s_{2},s_{3})$ are the side lengths with $s_{1}\geq s_{2}\geq s_{3}$. Since openable parts are typically plate-like, we define the front face $P_{\mathrm{front}}$ as the OBB face spanned by the two longest edges $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$ with unit normal $\mathbf{n}_{\mathrm{front}}=\mathbf{u}_{1}\times\mathbf{u}_{2}$. Let the initial articulation be $\phi(m)=(t,\mathbf{o},\mathbf{a},\rho)$. For prismatic joints, we update only the joint axis direction by aligning it with the front-face normal,
@@ III-E Interactive Scene Integration
Duplicate removal.  To ensure clean assembly of the final interactive scene, we perform a last deduplication pass. Let $\{S_{i}\}$ denote movable-part meshes with point sets $\{X_{i}\}$. We compute the IoU directly on these point sets, $\mathrm{IoU}(X_{i},X_{j})=\tfrac{|X_{i}\cap X_{j}|}{|X_{i}\cup X_{j}|}$. The procedure has two stages. Stage 1 (pairwise pruning): for every pair $(i,j)$, if $\mathrm{IoU}(X_{i},X_{j})\geq\tau_{\mathrm{dup}}$, we keep the mesh with greater information content (measured by point count) and discard the other, producing a reduced set $\{\hat{S}_{i}\}$ with point sets $\{\hat{X}_{i}\}$. Stage 2 (subdivision pruning): duplicates can also appear when a larger mesh can be represented by the union of several smaller, complementary meshes while no single pair exceeds $\tau_{\mathrm{dup}}$. We sort $\{\hat{S}_{i}\}$ by $|\hat{X}_{i}|$ in descending order and iterate from large to small so that larger parts can be explained by smaller ones first. For each $\hat{S}_{i}$, we collect candidate smaller meshes $\mathcal{C}_{i}=\{\,j:|\hat{X}_{j}|<|\hat{X}_{i}|,\ \mathrm{IoU}(\hat{X}_{i},\hat{X}_{j})\geq\tau_{\mathrm{low}}\,\}$, where $\tau_{\mathrm{low}}$ is a small IoU threshold, and enumerate their combinations. If there exists a subset $\mathcal{U}\subseteq\mathcal{C}_{i}$ such that $\mathrm{IoU}\!\left(\hat{X}_{i},\ \bigcup_{j\in\mathcal{U}}\hat{X}_{j}\right)\geq\tau_{\mathrm{dup}}$, we remove $\hat{S}_{i}$ as a redundant superset. In the end, we remove all redundant movable parts together with their associated inner boxes before assembling the final interactive scene.
@@
Scene assembly.  We integrate interactive objects into the original static scene by carving the background geometry using the world-frame OBBs of deduplicated movable parts. Let the original static scene mesh be $\mathbf{P}$ and the set of movable-part OBBs be $\{B_{j}\}$. To remove residual fragments from imperfect segmentation and spurious points in the original scan, we delete all vertices lying inside any $B_{j}$ together with their incident faces. This yields a cleaned background mesh $\mathbf{P}_{\mathrm{rem}}\;=\;\mathbf{P}\setminus\bigcup_{j}B_{j}$. We then assemble the final interactive scene by combining $\mathbf{P}_{\mathrm{rem}}$ with the interactive objects, ensuring a clean integration without interpenetration.
@@
Texture generation.  As both the original scene mesh and the segmented movable-part meshes provide only per-vertex colors, we generate texture maps based on InstantTexture [6]. We first perform UV unwrapping with Xatlas [29] on the movable parts and the cleaned remainder mesh. Given the UV parameterization, we rasterize each triangle in atlas space and compute per-texel colors by barycentric interpolation of the source vertex colors, yielding dense textures aligned to the unwrapped charts. We then repair uncovered texels caused by sampling gaps or occlusions via inpainting, followed by smoothing with low-pass filtering to attenuate aliasing across chart boundaries. Finally, the generated textures are bound to their corresponding meshes and, together with the refined articulation parameters, exported as .dae assets that are referenced in the scene description files.
@@
Export and simulator support.  To ensure portability across visualization and simulation stacks, we export each interactive scene in two widely adopted formats: URDF for the robotics ecosystem and USD for graphics/physics engines. The exports package the cleaned background, articulated interactive objects, and textures, enabling direct deployment in PyBullet [2], ROS [23], and Isaac Sim [20]. Fig. 5 shows the simulated interactive scenes generated by REACT3D from different inputs in Isaac Sim. For ROS, we additionally provide a lightweight GUI for manipulating the interactive scene during visualization and benchmarking (see Fig. 6).
@@
Method #Openable Objects $\tau^{*}=0.25$ $\tau^{*}=0.50$
@@
URDFormer 226 $0.429$ $0.226$ $0.296$ $0.361$ $0.190$ $0.249$
DRAWER (GT-mesh) 540 $0.434$ $0.141$ $0.213$ $0.326$ $0.106$ $0.159$
REACT3D 540 0.731 0.393 0.511 0.666 0.357 0.465